OpenStudy (anonymous):

solve the equation x/3 + x/4 = 1

myininaya (myininaya):

i would clear the fractions to make life easier

myininaya (myininaya):

do this by multiplying both sides by 12

myininaya (myininaya):

\[12(\frac{x}{3}+\frac{x}{4})=12(1)\]

myininaya (myininaya):

so what does that give us?

OpenStudy (anonymous):

3/3+3/4? 1.2

myininaya (myininaya):

12/3=4 12/4=3 12(1)=12 12x/3=4x 12x/4=3x right? so our equation is 4x+3x=12 7x=12 What is the final operation we must performed in order to isolate x?

OpenStudy (anonymous):

idk how?

OpenStudy (anonymous):

?

myininaya (myininaya):

To undo multiplication we need to ________

OpenStudy (anonymous):

divide

myininaya (myininaya):

yep

myininaya (myininaya):

7x=12 we divide both sides by ?